# Rethinking Creativity Evaluation: A Critical Analysis of Existing Creativity Evaluations

### v2

@@ Abstract
We systematically examine, analyze, and compare representative creativity measures—creativity index, perplexity, syntactic templates, and LLM-as-a-Judge—across diverse creative domains, including creative writing, unconventional problem-solving, and research ideation. Our analyses reveal that these metrics exhibit limited consistency, capturing different dimensions of creativity. We highlight key limitations, including the creativity index’s focus on lexical diversity, perplexity’s sensitivity to model confidence, and syntactic templates’ inability to capture conceptual creativity. Additionally, LLM-as-a-Judge shows instability and bias. Our findings underscore the need for more robust, generalizable evaluation frameworks that better align with human judgments of creativity. We release the datasets and evaluation code: https://github.com/lichun-19/creative_eval
@@ 1 Introduction
Creativity is essential to human progress across art, daily problem-solving, and scientific discovery (cropley2006role; feist1998meta; simonton2004creativity). Large language models (LLMs) perform well on structured tasks such as question answering and logical reasoning (goel2023; cabral2024; pu2023summarizationalmostdead; li2024flexkbqaflexiblellmpoweredframework), but their creative abilities remain limited (Sawicki2023OnTP; sawicki2023bitsgrassdoesgpt; chakrabarty2024; mirowski). As interest in machine creativity grows, so too does the need for reliable evaluation metrics that can effectively benchmark LLM performance across creative domains.
@@
The inherently multidimensional and domain-specific nature of creativity has led to a diverse range of evaluation methods. Prior work has proposed various automatic metrics to capture different aspects of creativity, such as semantic distance to measure novelty in text generation (zhao2019; haque2022), and metrics like the creativity index and syntactic templates for assessing linguistic creativity (li2016; lu2025salieri; liu2024infinigramscalingunboundedngram; shaib2024). LLMs have also been used as evaluators in tasks like the Alternative Uses Task (AUT) and creative writing (Goes2023; organisciak; distefano2024; luchini2025automated; luchini2025multilingual). These methods highlight the diversity of creativity and the complexity of evaluating creativity. This raises an important question: how can we select evaluation metrics accurately reflecting human-perceived creativity across various domains?
@@
To address this question, we analyze four representative creativity metrics, each reflecting a different level of granularity in creativity evaluation. W use the creativity index to assess phrase-level creativity, perplexity to capture token-level diversity, syntactic templates to evaluate structural creativity, and LLM-as-a-Judge for a holistic assessment (merrill2024evaluating; crossley2016idea; nguyen2024). To evaluate the alignment between these metrics and human perceptions of creativity, we apply them across three domains where machine creativity has been extensively studied (ismayilzada2024creativityaiprogresseschallenges): artistic creativity (e.g., creative writing), unconventional problem-solving, and research ideation. For each domain, we curate a dataset of both creative and uncreative examples, reflecting human judgments, and assess each metric’s performance within and across these domains.
@@
Our analysis reveals that the four creativity metrics exhibit limited consistency with each other. Perplexity and syntactic templates, in particular, show inconsistent performance across domains, reflecting their sensitivity to task-specific language patterns and limitations in generalizability. We further analyze each metric’s behavior:
  • Creativity Index mainly reflects lexical diversity, with its reliance on the chosen $n$-gram range affecting its reliability.
  • Perplexity, while associated with diversity and fluency, is heavily influenced by model confidence, making it a poor indicator of creativity in contexts that prioritize novelty.
  • Syntactic Templates can detect structural patterns but fail to capture conceptual creativity, especially in unconventional solutions expressed through conventional language.
  • LLM-as-a-Judge tends to produce biased predictions and is unstable across different prompts, raising concerns about its reliability as an autonomous evaluator.
@@ 2 Background on creativity evaluation
As LLMs advance in generating creative content, the need for effective evaluation methods has grown. In the following, we review existing metrics and discuss their applicability to evaluating machine-generated content.
@@
Classical psychological tests such as the Torrance Tests of Creative Thinking (torrance1966) (assessing fluency, flexibility, originality, and elaboration), the Divergent Association Task (olson2021) (semantic distance and associative thinking), the Remote Associates Test (mednick1962associative) (linking seemingly unrelated words), and the Alternative Uses Task (guilford1967) (generating novel uses for common objects) capture different facets of creative thinking. While these tests are well-established for assessing human creativity, their suitability for evaluating and optimizing machine learning models, such as LLMs, is limited. First, traditional metrics, such as fluency and flexibility, become less informative as LLMs can generate extensive content with ease (guzik2023; lu2024llm). Second, these tests require significant human effort and cost, posing challenges for integration into model training loops (torrance1966; olson2021; guilford1967). Finally, the reliance on subjective human evaluation could introduce potential biases (beaty2021).
@@
As large language models are increasingly applied to creative domains such as story writing (fan2018; akoury2020; gomez2023) and music composition (ding2024songcomposerlargelanguagemodel; parker2024; yuan2024chatmusician), the scope of evaluation has expanded as well. Several benchmarks have been introduced to assess the creative capacity of LLMs (lu2024benchmarking; zhang2025noveltybench; atmakuru2024cs4). These benchmarks highlight a system-level view of creativity and provide useful comparisons across different models. Our focus, however, is on the level of individual outputs: developing automatic metrics that can assess creativity in the open-ended settings where it often naturally arises, beyond the constraints of predefined tasks.
@@
For this output-level focus, several automated creativity measures, including linguistic diversity (li2016; lu2025salieri; liu2024infinigramscalingunboundedngram; shaib2024), text perplexity (fan2018), and LLM-based judgment (ruan2025; ye2024; lin2024; pai2025billy), have been proposed as scalable alternatives to human assessment. However, LLM-based evaluation is susceptible to biases, such as positional bias and preference for model-generated outputs (shi2024; panickssery2024; ye2024bias), raising concerns about its reliability, and few works have examined whether these metrics meaningfully capture conceptual novelty: a gap we address in this study.
@@
To close this gap, our framework analyzes creativity through a technical segmentation of outputs at the token, phrase, and overall levels, enabling scalable, fine-grained evaluation that reduces human bias and cost. Complementing prior work that draws on human judgments, through adaptations of psychological tests (chakrabarty2024; tian2024a; lu2024llm) or qualitative frameworks emphasizing dimensions such as surprise, diversity, and novelty (ismayilzada2024evaluating; ismayilzada2025creative), our approach provides an interpretable foundation for systematic, automated analysis. Together, these perspectives capture the full dimensions of creativity.
In this work, we aim to systematically examine, analyze, and compare existing creativity measures across diverse creative domains. By offering a rigorous analysis, we seek to inform future research and guide the development of more effective creativity assessment methods.
@@ 3 Core domains of creativity selected for evaluation
Our goal is to systematically analyze and compare existing creativity measures across diverse creative domains. Section 3.1 introduces the domains we selected for evaluation, and Section 3.2 describes the datasets we curated for these domains.
@@ 3.1 Domains
Creativity is an incredibly broad concept spanning multiple disciplines; to aid the evaluation of existing creativity metrics, we first identify the core domains of creativity. We follow the broad categorization by ismayilzada2024creativityaiprogresseschallenges and select three domains: creative writing, unconventional problem-solving, and research ideation.
Each domain offers a unique aspect of creativity: creative writing uses artistic and aesthetic approaches, unconventional problem-solving uses functional thinking, and research ideation uses logical synthesis of existing facts.
Creative writing. We focus on creative writing (Creative writing) as a representative subfield for artistic creativity. It involves generating original, aesthetically engaging content that elicits an emotional or intellectual response in audiences and plays a vital role in individual well-being and social connection (jean2024creative; kaimal2017functional; fancourt2019how; kim2015community; felton2015widening).
Problem-solving. Unconventional problem-solving (Problem-solving) involves identifying creative yet effective solutions to unconventional problems. It often involves breaking away from conventional approaches, reframing problems from different perspectives, and overcoming cognitive biases such as functional fixedness. Effective problem solvers leverage a combination of domain knowledge, lateral thinking, and adaptability to generate innovative and practical outcomes. This domain of creativity requires both divergent and convergent thinking, and is widely studied in chen2024boosting.
Research ideation. Research ideation (Research ideation) involves generating, exploring, and evaluating novel scientific ideas that can lead to advancements in knowledge and practice (lu2024ai). This domain is marked by the ability to challenge existing paradigms, synthesize seemingly unrelated concepts across disciplines, and propose innovative hypotheses that push the boundaries of current understanding. Researchers often rely on both divergent thinking to generate a wide range of possibilities and convergent thinking to refine and evaluate the feasibility of these ideas. This creative process is crucial for breakthroughs in science and technology, driving forward the discovery of new theories, methods, and applications that can address complex global challenges.
@@
Problem Solving Finding creative solutions to unconventional problems. Given items {A, B, C}, how can we achieve {purpose} when these items weren’t designed for that purpose?
@@ 3.2 Datasets
To investigate the performance of existing creativity metrics across domains, we curate a dataset for each domain, with examples described in Table 1. Each dataset contains two sets of samples (creative and uncreative) labeled based on human perception of creativity. Section H presents representative creative and uncreative examples for each domain. To validate reliability, we conduct a pairwise verification task in which annotators identify the more creative example from 30 random pairs per dataset, achieving 73% agreement with the human-labeled creative samples. We provide the details of annotators and instructions in Section D.
For Creative writing, we build upon datasets from a prior work (tian2024a), which showed LLM-generated narratives were less diverse in story arcs and had worse pacing compared to human narratives; we therefore label the dataset’s 150 LLM-generated narratives as uncreative and its 150 human narratives as creative, with supporting details in Section C. We augment the dataset with human-written synopses scraped from movie summary sites MoviePooper and The Movie Spoiler.
For Problem-solving, we adapt the MacGyver dataset for functional fixedness (tian2024b), which contains a collection of problem-solving scenarios designed to assess the ability to think beyond conventional uses of everyday objects. We filter $573$ unconventional solutions and extract the tools mentioned in each problem to maintain similar settings across examples. We then augment the dataset by prompting LLMs to generate corresponding conventional problem-solution pairs, the validity of which we verify through human annotation. Both unconventional and conventional solutions were carefully rephrased to ensure consistent expression and style across the dataset.
For Research ideation, we collect data from OpenReview submissions to ICLR 2024, which has published soundness, contribution, and presentation scores for all accepted and rejected submissions. To control for variables other than creativity in research ideas, we group the papers according to their soundness and presentation scores. Section A details the curation process, which enables us to focus on the creativity of papers as the primary differentiating factor between accepted and rejected papers as implemented by zhao2025review. We balance our sample size across groups, resulting in a final dataset of $513$ creative/accepted papers and $513$ uncreative/rejected papers.
@@ 4 Creativity metrics selected for evaluation
Creativity Index (CI) quantifies textual originality by measuring overlap with web corpora with the DJ Search algorithm (lu2025salieri). It estimates $L$-uniqueness, the proportion of words appearing in novel $n$-gram contexts of length $L$ or greater, and averages these scores across a range of $L$ values. In this study, we focus on exact verbatim matches and implement CI using the Infini-Gram algorithm (liu2024infinigramscalingunboundedngram). See Section B for details.
Perplexity (PPL) measures the statistical unexpectedness of text based on probability distributions (jelinek1977). It indicates how well a probability model predicts a sample, where lower probability is higher perplexity and therefore the content is less expected and possibly more creative.
Syntactic templates examine structural diversity in text by identifying commonly-used part-of-speech patterns (e.g., [DET NOUN VERB DET ADJ NOUN]) (shaib2024). It uses three key metrics:
@@
LLM-as-a-Judge uses LLMs to assess creativity at the conceptual level, complementing the linguistic metrics above. Building on recent work highlighting LLMs’ promise for creativity assessment (organisciak; distefano2024; luchini2025automated; luchini2025multilingual), we conduct extensive ablations across various dimensions, e.g., prompt sensitivity, model bias, model consistency, and potential data contamination, as described in Section F, and report the results of the best-performing configuration. Our approach employs a multi-pronged evaluation framework that targets idea-level creativity through chain-of-thought prompting, aspect-based evaluation, and clearly defined scoring criteria.
@@ 5.1 Quantitative results
Problem Solving Unconventional 0.76 $\pm$ 0.08 13.98 $\pm$ 4.65 6.158 38.7% (0.039) 3.02 $\pm$ 0.87
@@
Creativity Index (CI). CI effectively captures the differences in Creative writing; however, it yields comparable scores between unconventional and conventional solutions in Problem-solving, as well as between accepted and rejected paper introductions in scientific ideation. As discussed in Section 5.2.1, prior work often uses unique n-grams to evaluate novelty (crossley2016idea; merrill2024evaluating). However, our results suggest CI reflects primarily lexical diversity which is more relevant in the writing domain rather than originality of ideas, limiting its effectiveness in capturing conceptual creativity. Furthermore, Section 5.2.2 highlights additional factors, such as the choice of the n-gram range (L uniqueness), which can significantly influence the CI scores and ultimately undermine the reliability of reflecting the linguistic creativity itself. Together with its dependence on the reference corpus, this strong sensitivity limits CI’s generality across domains and evaluation setups.
@@
Although some studies have proposed PPL as an indicator of surprise, diversity, or creativity (basu2021mirostat), it was originally developed as a fluency metric and is heavily shaped by trade-offs between model confidence and output smoothness (guo2024curiousdeclinelinguisticdiversity). Our findings show the score distributions for creative and non-creative texts overlap substantially across all domains, offering little discriminative signal. We therefore suggest PPL remains useful for monitoring fluency or guiding fine-tuning (guo2024curiousdeclinelinguisticdiversity; wu2025mitigating), but should not be treated as a reliable metric of creativity.

### v3

@@ Abstract
We examine, analyze, and compare four representative creativity measures—perplexity, LLM-as-a-Judge, the Creativity Index (CI; measuring n-gram overlap with web corpora), and syntactic templates (detecting repetition of common part-of-speech patterns)—across the diverse creative domains, such as creative writing, unconventional problem-solving, and research ideation. For each domain, we compile datasets with human-aligned creative and uncreative examples and evaluate each metric’s ability to discriminate between the two sets. Our analyses reveal limited consistency both across domains and metrics, as metrics that distinguish creativity in one domain fail in others (e.g., CI correctly distinguishes in creative writing but fails in problem-solving), and different metrics often disagree on the same data points (e.g., CI suggests one set to be more creative, while perplexity indicates the other set to be more creative.) We highlight key limitations, such as perplexity reflecting fluency rather than novelty; LLM-as-a-Judge producing inconsistent judgments under minor prompt variations and exhibiting bias towards particular labels; CI primarily measuring lexical diversity, with high sensitivity to implementation choices; and syntactic templates being ineffective in settings dominated by formulaic language. Our findings underscore the need for more robust, generalizable evaluation frameworks that better align with human judgments of creativity. We release the datasets and evaluation code: https://github.com/lichun-19/creative_eval.
@@ 1 Introduction
Creativity is essential to human progress across art, daily problem-solving, and scientific discovery (Cropley, 2006; Feist, 1998; Simonton, 2004). Large language models (LLMs) perform well on structured tasks such as question answering and logical reasoning (Goel et al., 2023; Cabral et al., 2024; Pu et al., 2023; Li et al., 2024), but their creative abilities remain limited (Sawicki et al., 2023a, b; Chakrabarty et al., 2024; Mirowski et al., 2023). As interest in machine creativity grows, so too does the need for reliable evaluation metrics that can effectively benchmark LLM performance across creative domains.
@@
The inherently multidimensional and domain-specific nature of creativity has led to a diverse range of evaluation methods. Prior work has proposed various automatic metrics to capture different aspects of creativity, such as semantic distance to measure novelty in text generation (Zhao et al., 2019; Haque et al., 2022), and metrics like the Creativity Index and syntactic templates for assessing linguistic creativity (Li et al., 2016; Lu et al., 2024c; Liu et al., ; Shaib et al., 2024). LLMs have also been used as evaluators in tasks like the Alternative Uses Task (AUT) and creative writing (Góes et al., 2023; Organisciak et al., 2023; Distefano et al., 2024; Luchini et al., 2025a, b). These methods highlight the diversity of creativity and the complexity of its evaluation. This raises an important question: how can we select evaluation metrics accurately reflecting human-perceived creativity across domains?
@@
To address this question, we analyze four representative creativity metrics, each reflecting a different level of granularity in creativity evaluation as illustrated in Figure 1. We adopt the commonly used definition of creativity (Guerberof-Arenas and Toral, 2022; Mukherjee and Chang, 2023; Ismayilzada et al., 2025), which treats creativity as the combination of novelty and acceptability (also referred to as quality or usefulness, i.e., the idea is appropriate or fit for its purpose and constraints). In our setting, we emphasize novelty, as our dataset construction process already filters out infeasible or low-quality examples. We use the Creativity Index (Lu et al., 2024c) to assess phrase-level creativity by computing n-gram overlap with large-scale web corpora, perplexity to capture token-level diversity, syntactic templates (Shaib et al., 2024) to evaluate structural creativity, and LLM-as-a-Judge for a holistic assessment (Merrill et al., 2024; Crossley et al., 2016; Nguyen, 2024). To evaluate the alignment between these metrics and human perceptions of creativity, we apply them across three domains where machine creativity has been extensively studied (Ismayilzada et al., 2024a): artistic creativity (e.g., creative writing), unconventional problem-solving, and research ideation. For each domain, we curate a dataset of both creative and uncreative examples, reflecting human judgments, and assess each metric’s performance within and across these domains.
@@
Our analysis reveals that the four creativity metrics exhibit limited consistency both across domains and with one another. We further analyze each metric’s behavior:
  • Creativity Index is weakened by its sensitivity to the choice of L-uniqueness and the selection of reference corpora, e.g., difference between two cases can vary by a factor of two when using L = 5–7 versus L = 5–11.
  • Perplexity, while associated with diversity and surprise (Basu et al., 2021), exhibits substantial overlap in creative and uncreative score distributions across all domains.
  • Syntactic templates are weak in settings where novel ideas are often expressed through formulaic language, such as research ideation and problem-solving.
  • LLM-as-a-Judge is highly prompt-sensitive and prone to one-sided bias across models; furthermore, repeated evaluations of the same instance by the same judge exhibit only 40% consistency across three runs.
@@ 2 Background on creativity evaluation
As LLMs advance in generating creative content, the need for effective evaluation methods has grown. We review existing metrics and discuss their applicability to evaluating machine-generated content.
@@
Classical psychological tests, including the Torrance Tests of Creative Thinking (Torrance, 1966), Divergent Association Task (Olson et al., 2021), Remote Associates Test (Mednick, 1962), and Alternative Uses Task (Guilford, 1967), capture different facets of creative thinking, but are limited for evaluating LLMs. Firstly, traditional metrics like fluency become less informative as LLMs generate extensive content with ease (Guzik et al., 2023; Lu et al., 2024b); secondly, these tests require significant human effort (Torrance, 1966; Olson et al., 2021; Guilford, 1967); and finally, subjective evaluation introduces potential biases (Beaty and Johnson, 2021).
@@
As LLMs are increasingly applied to creative domains such as story writing (Fan et al., 2018; Akoury et al., 2020; Gómez-Rodríguez and Williams, 2023) and music composition (Ding et al., 2024; Parker et al., 2024; Yuan et al., 2024), several benchmarks have been introduced to assess their creative capacity (Lu et al., 2024d; Zhang et al., 2025; Atmakuru et al., 2024). These benchmarks provide system-level comparisons across models. Our focus, however, is on the level of individual outputs: developing automatic metrics that can assess creativity in open-ended settings beyond the constraints of predefined tasks.
@@
Automated measures including linguistic diversity (Li et al., 2016; Lu et al., 2024c; Liu et al., ; Shaib et al., 2024), text perplexity (Fan et al., 2018), and LLM-based judgment (Ruan et al., 2024; Ye et al., 2025; Lin et al., 2024; Pai et al., 2026) have been proposed as scalable alternatives to human assessment. However, LLM-based evaluation is susceptible to biases such as positional bias and preference for model-generated outputs (Shi et al., 2024; Panickssery et al., 2024; Ye et al., 2024), and few works have examined whether these metrics meaningfully capture conceptual novelty, which is a gap we address in this study.
@@
To close this gap, our framework analyzes creativity through a technical segmentation at the token, phrase, and overall levels, complementing prior work that draws on human judgments using adapted psychological tests (Chakrabarty et al., 2024; Tian et al., 2024a; Lu et al., 2024b) or qualitative frameworks emphasizing surprise, diversity, and novelty (Ismayilzada et al., 2024b, 2025). By systematically examining existing measures across diverse creative domains, we provide an interpretable foundation for systematic, automated analysis and guide the development of more effective assessment methods for future research.
@@ 3 Core domains of creativity selected for evaluation
To systematically analyze existing creativity measures across diverse creative domains, we introduce the domains we selected in Section 3.1, and describe the curation process in  Section 3.2.
@@ 3.1 Domains
Creativity is an incredibly broad concept spanning multiple disciplines; to aid the evaluation of existing creativity metrics, we first identify the core domains of creativity.
We follow the broad categorization by Ismayilzada et al. (2024a) and select three domains representing distinct aspects of creativity: creative writing (artistic and aesthetic approaches), unconventional problem-solving (functional thinking), and research ideation (logical synthesis of existing facts).
Creative writing. Creative writing serves as a representative sub- field for artistic creativity, which involves generating original, aesthetically engaging content that elicits emotional or intellectual responses (Jean-Berluche, 2024; Kaimal et al., 2017; Fancourt et al., 2019; Kim, 2015; Felton et al., 2015).
Problem-solving. Problem-solving involves identifying creative yet effective solutions by breaking from conventional approaches and overcoming cognitive biases such as functional fixedness, with the adaptability to generate innovative and practical outcomes. This domain of creativity requires both divergent and convergent thinking (Chen et al., 2024).
Research ideation. Research ideation involves generating and evaluating novel scientific ideas by challenging existing paradigms, synthesizing concepts across disciplines (Lu et al., 2024a). This creative process is crucial for driving breakthroughs that advance the discovery of new theories, methods, and applications capable of addressing complex global challenges.
@@
Problem-solving Finding creative solutions to unconventional problems. Given items {A, B, C}, how can we achieve {purpose} when these items weren’t designed for that purpose?
@@ 3.2 Datasets
To investigate the performance of existing creativity metrics across domains, we curate a dataset for each domain, with examples described in Table 1. Each dataset contains two sets of samples (creative and uncreative) labeled based on human perception of creativity. Section H presents representative creative and uncreative examples for each domain. To validate reliability, we conduct a pairwise verification task in which annotators identify the more creative example from 30 random pairs per dataset, achieving 73% agreement with the human-labeled creative samples. Annotation was conducted by six graduate students with professional English proficiency (CEFR level $\geq$ B2). Each data point was independently annotated by two annotators, and inter-annotator agreement was subsequently computed on the paired annotations. Annotation instructions are provided in Section D.
For Creative writing, we build upon datasets from Tian et al. (2024a), which showed LLM-generated narratives were less diverse in story arcs and had worse pacing compared to human narratives; we therefore label the dataset’s 150 LLM-generated narratives as uncreative and its 150 human narratives as creative, with supporting details in Section C. We augment the dataset with human-written synopses scraped from movie summary sites MoviePooper and The Movie Spoiler.
For Problem-solving, we adapt the MacGyver dataset for functional fixedness (Tian et al., 2024b), which contains a collection of problem-solving scenarios designed to assess the ability to think beyond conventional uses of everyday objects. We filter $573$ unconventional solutions and extract the tools mentioned in each problem to maintain similar settings across examples. We then augment the dataset by prompting LLMs to generate corresponding conventional problem-solution pairs, the validity of which we verify through human annotation. Both unconventional and conventional solutions were carefully rephrased to ensure consistent expression and style across the dataset.
For Research ideation, we collect data from OpenReview submissions to ICLR 2024, which has published soundness, contribution, and presentation scores for all accepted and rejected submissions. To control for variables other than creativity in research ideas, we group the papers according to their soundness and presentation scores. Section A details the curation process, which enables us to focus on the creativity of papers as the primary differentiating factor between accepted and rejected papers as implemented by Zhao and Zhang (2025). We balance our sample size across groups, resulting in a final dataset of $513$ creative/accepted papers and $513$ uncreative/rejected papers.
@@ 4 Creativity metrics selected for evaluation
Creativity Index (CI) quantifies textual originality by measuring overlap with web corpora with the DJ Search algorithm (Lu et al., 2024c). It estimates $L$-uniqueness, the proportion of words appearing in novel $n$-gram contexts of length $L$ or greater, and averages these scores across a range of $L$ values. In this study, we focus on exact verbatim matches and implement CI using the Infini-Gram algorithm (Liu et al., ). Section B provides additional details and shows that near-verbatim matching yields results similar to exact matching.
Perplexity (PPL) measures the statistical unexpectedness of text based on probability distributions (Jelinek et al., 1977). It indicates how well a probability model predicts a sample, where lower probability is higher perplexity and therefore the content is less expected and possibly more creative.
Syntactic templates examine structural diversity in text by identifying commonly-used part-of-speech patterns (e.g., [DET NOUN VERB DET ADJ NOUN]) (Shaib et al., 2024). It uses three key metrics:
@@
LLM-as-a-Judge uses LLMs to assess creativity at the conceptual level, complementing the linguistic metrics above. Building on recent work highlighting LLMs’ promise for creativity assessment (Organisciak et al., 2023; Distefano et al., 2024; Luchini et al., 2025a, b), we conduct extensive ablations across various dimensions, e.g., prompt sensitivity, model bias, model consistency, and potential data contamination, as described in Section F, and report the results of the best-performing configuration. Our approach employs a multi-pronged evaluation framework that targets idea-level creativity through chain-of-thought prompting, aspect-based evaluation, and clearly defined scoring criteria.
@@ 5.1 Quantitative results
Problem-solving unconventional 0.76 $\pm$ 0.08 13.98 $\pm$ 4.65 6.158 38.7% (0.039) 3.02 $\pm$ 0.87
@@
Creativity Index (CI). CI effectively captures the differences in Creative writing; however, it yields comparable scores between unconventional and conventional solutions in Problem-solving, as well as between accepted and rejected paper introductions in scientific ideation. As discussed in Section 5.2.1, prior work often uses unique n-grams to evaluate novelty (Crossley et al., 2016; Merrill et al., 2024). However, our results suggest CI reflects primarily lexical diversity which is more relevant in the writing domain rather than originality of ideas, limiting its effectiveness in capturing conceptual creativity. Furthermore, Section 5.2.2 highlights additional factors, such as the choice of the n-gram range (L uniqueness), which can significantly influence the CI scores and ultimately undermine the reliability of reflecting the linguistic creativity itself. Together with its dependence on the reference corpus, this strong sensitivity limits CI’s generality across domains and evaluation setups.
@@
Although some studies have proposed PPL as an indicator of surprise, diversity, or creativity (Basu et al., 2021), it was originally developed as a fluency metric and is heavily shaped by trade-offs between model confidence and output smoothness (Guo et al., 2024). Our findings show the score distributions for creative and uncreative texts overlap substantially across all domains, offering little discriminative signal. We therefore suggest PPL remains useful for monitoring fluency or guiding fine-tuning (Guo et al., 2024; Wu et al., 2025), but should not be treated as a reliable metric of creativity.